Ticket Number: HEA-NOC/19990129-01 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: Avoca.heanet.ie
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 19990201 18:00 UTC
Ticket Closed: 19990201 20:31:49 UTC Problem End: 19990201 20:25
Avoca.heanet.ie was brought down at 18:00 on Monday, Feb 1st
to patch & reconfigure the kernel.
The following services were be offline:
www.heanet.ie (ftp.heanet.ie, www.hea.net)
listserv.heanet.ie (mail will queue)
cache.heanet.ie (ICP peers will bypass transparently)
Install patch kit, rebuild kernel, reboot.
Further configuration & reboot may be necessary.
19990201 20:30 UTC davew
Patch went ok although many patches could not be installed because of
system changes in the past. System was brought back at 19:00 and kernel
changes began. These caused severe trouble and multiple reboots were
necessary to get the system stable again. The problem changes have been
identified & will follow up with DEC for more information.
All services are back; LISTSERV backlogs are now clearing.
Time to Fix: